one number exceeds another number by 6 when twice the smaller number is added to three times the smaller number,the result is 32.find the two numbers
Answer by ptaylor(1894)

one number exceeds another number by 6 when twice the smaller number is added to three times the smaller number,the result is 32.find the two numbers
SHOULDN'T THAT BE "----TWICE THE LARGER NUMBER IS ADDED TO----"
Let x=smaller number
Then x+6=larger number
2(x+6)=twice the larger number
3x=three times the smaller number
Now we are told that:
2(x+6)+3x=32 clear parens
2x+12+3x=32 subtract 12 from both sides
2x+3x+12-12=32-12
5x=20
x=4---------------smaller number
x+6=4+6=10-----------larger number
CK
Larger number exceeds the smaller number by 6-------YES
twice the larger number (20) plus three times the smaller number (12) equals 32
YES
Hope this helps----ptaylor